# The multimessenger view of Pulsar Timing Array black holes with the Horizon-AGN simulation

### v1

@@ 6 Properties of MBHs and host galaxies
Based on the analysis presented in the previous sections, we focus here on binaries with chirp mass $10^{8.35}\,{\rm M_{\odot}}$ and $0.05<z<1.2$ as the dominant component of the background (Section 4.2.2) and the selected individual CW candidates in population modeling for circular binaries (Section 5.2.2) as representative of the population. As a general point of information, a given MBH may appear in the catalog multiple times with multiple redshifts as it merges with different MBHs. This is the case for the most massive MBHs, hosted in very massive galaxies, since they experience a large number of mergers in their lifetime. Physically, this does not mean that a given binary in the Universe would be detected multiple times. However, if a simulation is a realistic representation (of a finite volume) of our Universe, the MBH binaries populating a given redshift layer will have similar properties as those merging at the corresponding redshift in the simulation. For a sufficiently large simulation box, which probes the large scale structure of the Universe, this is a reasonable assumption.
@@
What makes an MBHB a good CW candidate physically? Besides its position in the sky, near pulsars, and its inclination, which we assumed to be both random in our analysis, the only physical quantities from the simulation are chirp mass and redshift. We consider binaries that make the best CW candidates as those that appear most often in the various universe realizations. The best CW candidates are not necessarily those that have the largest SNR. The largest SNR are obtained for the largest chirp masses and the lowest redshifts (luminosity distance). The largest chirp masses have however the shortest residence times in the PTA band (Mingarelli et al., 2017), making it harder to catch them in the act. The lowest redshift sources are rarer because of the small volume probed. As a result, the binaries associated with the very high SNR are also appearing very rarely in the universe realizations. We therefore consider as ”high probability” CW candidates those that appear multiple times - at least 5 times in this paper - across our 2 0002\,000 realizations.
@@
In the following we will consider both the simulation box, which is an evolving patch of the Universe that we can analyze at different times, and the simulation lightcone, which mimics what and how an observational survey would “see” the simulation. When we move from the simulation box to the lightcone generated for the simulation (Laigle et al., 2019) a given MBH appears only once, and the cosmic time at which the MBH appears in the lightcone may or not coincide with the time of a MBH merger. This means that only a small fraction of merging binaries are contained in the lightcone.
@@
In Figure 9, we compare the properties of mergers dominating the background in the simulation box which has a fixed co-moving volume (teal), in the simulation lightcone (light green) which simulates the source distribution seen by an observer, and the CW candidates (red). In these distributions, we include CW candidates accounting for their multiplicity across universe realizations, as the multiplicity of a given candidate in the various realizations provides information on how strong a candidate is as it provides information on its likelihood. In a cosmological volume, and in a lightcone that has increasing volume as redshift increases, the background is dominated by the more abundant MBHs with mass $\sim 10^{8.5}\,{\rm M_{\odot}}$, rather than the stronger but rarer GW sources with larger mass. CW candidates are instead mainly found at low redshift and have higher MBH masses; some can have also relatively low masses, down to $\sim 3\times 10^{7}\,{\rm M_{\odot}}$, but they are all at $z<0.1$. The redshift distribution extends to higher redshifts in the lightcone, because of the larger volume probed with increasing redshift, while it is limited to lower redshift for CW candidates, because the strength of the GW signal decreases with increasing luminosity distance, hence redshift. The degeneracy between luminosity distance (redshift) and chirp mass (see Equations 22 and 30) allows the presence of CW candidates up to $z=0.4$, provided that they have chirp mass $\mathcal{M}_{\rm c}\sim 10^{9.5}\,\,{\rm M_{\odot}}$. The distribution of the masses of the host galaxies reflects the $\mathcal{M}_{\rm c}$ and $z$ distributions for the background, while for CW candidates it is dominated by massive galaxies, in line with the larger MBH masses with respect to the background. These results are in agreement with previous and concurrent investigations, e.g., Sesana et al. (2009); Cella et al. (2025); Truant et al. (2025).
@@
We examine the relation between binaries and environment in Figure 10. Here, $\mathcal{M}_{\rm c}$ is shown as a function of the halo mass of the host. We note some “undermassive binaries”, with very low chirp mass with respect to the host galaxy/halo: these are cases where the mass ratio in the binary is very small, and therefore the chirp mass differs significantly from the binary mass, which is what scales with host properties in the simulation (shown in Volonteri et al., 2020). The binaries dominating the background in the simulation box are hosted mainly in galaxies and halos only relatively more massive than the Milky Way, while CW candidates are for the most part hosted in group-like main halos (red squares) with $M_{\rm halo}>10^{13}\,{\rm M_{\odot}}$. CW candidates that are hosted in lower mass halos are all in subhalos (orange crosses), meaning that the subhalo has been affected by stripping and it is part of a larger halo [^10] The plume of binaries in the background with halo mass $10^{12}\,{\rm M_{\odot}}$ is also caused by subhalos.; the high chirp mass with respect to halo mass is explained by stripping of the halo, which has therefore lost part of its mass (see the discussion in Volonteri et al., 2016, for the general MBH population). This implies that CW candidates are best looked for in galaxy groups and clusters, even more so than for the general population of MBH binaries (Rosado & Sesana, 2014; Izquierdo-Villalba et al., 2023; Saeedzadeh et al., 2024). Using groups/clusters as priors for where to look for CW candidates in the Bayesian search of PTA data could be therefore a possible strategy.
@@
Furthermore, the MBHs in the central galaxies of overdensities experience a large number of mergers, up to $\sim 20-25$ for the MBHBs that dominate the nHz background. This is a manifestation of the significant fraction of MBH mass gained through MBH mergers for these galaxies (Volonteri & Ciotti, 2013; Dubois et al., 2014b; Kulier et al., 2015). This occurs because of a combination of enhanced galaxy mergers and decreased gas availability for accretion for high-mass MBHs hosted in massive central galaxies of groups and clusters.
@@
In Figure 11 we compare the Eddington ratios and star formation rate (SFR) of background sources in the box vs CW candidates. We average SFR over 100 Myr since most star formation diagnostics probe SFR on such timescales and we average $f_{\rm Edd}$ over 1 Myr to smooth the variability. Also in this case we include CW candidates accounting for their multiplicity across universe realizations. The Eddington ratio of the binaries is generally low, $10^{-2}$, typical for MBHs of this mass and redshift in the simulation as well as in observations (Volonteri et al., 2016). Accretion is even lower for CW candidates, since they are at very low redshift, in very massive galaxies. The same considerations apply to SFR, in agreement with previous results which find that the host have low SFR or specific SFR (Saeedzadeh et al., 2024; Truant et al., 2025; Cella et al., 2025). Since we have considered delayed mergers rather than numerical mergers, the time elapsed from galaxy merger to binary observation is sufficiently long that any galaxy merger-driven increase in accretion rate or SFR has decayed (Dong-Páez et al., 2023).
@@
Examples of hosts of CW candidates in false gri colors are shown in Figure 12 and Figure 13. The MBHB is located at the center of the image. The first three images (Figure 12 and Figure 13 left) show two low-redshift cases consistently appearing in the realizations, while Figure 13, right, shows a rarer high-redshift ($z=0.42$) case. The images highlight that CW candidates are preferentially found in the center of galaxy groups and clusters, in agreement with the results shown in Figure 10.
@@ 7 Electromagnetic emission from galaxies and MBHs
Figure 14: Electromagnetic properties of MBHBs and their host galaxies for CW candidates. The left panels show B, V, and K magnitudes; the right panels X-ray flux at [2-10] keV, radio flux at 2 GHz and bolometric luminosity (the horizontal line marks $L_{\rm bol}=10^{45}\,\rm{erg}\,{\rm s^{-1}}$), as a typical value for quasar luminosities. For MBHs with $f_{\rm Edd}<10^{-3}$ we show an optimistic upper limit to the brightness ignoring the correction to radiative efficiency.
@@ 7.1 Methodology
Galaxies magnitudes are computed as described in Laigle et al. (2019). In brief, we compute the galaxy rest-frame spectral energy distribution by summing up the contribution of all stellar particles (knowing their ages and metallicities), assumed to behave as single stellar populations. We use the stellar population synthesis model of Bruzual & Charlot (2003) assuming a Chabrier (2003) initial mass function. The dust column density and optical depth along the line of sight of each star particle is computed assuming that 30% of metals in the galaxy gas distribution are locked into dust grains. We use the $R_{\rm V}=3.1$ Milky Way dust grain model by Weingartner & Draine (2001). Apparent total magnitudes are computed on the redshifted spectra, using the $\rm B$, $\rm V$ and $\rm K_{\rm s}$ filters.
In X-rays, the galaxy emission is dominated by X-ray binaries composed of a compact object accreting from a companion star. Empirical models that allow one to estimate the combined effect of low-mass X-ray binaries, where the companion is a low-mass star and the emission depends on the total galaxy mass, and high-mass X-ray binaries, where the companion is a high-mass star and the emission depends on the SFR. These models are based either on individual observations of nearby galaxies (e.g. $50\rm\,Mpc$, Lehmer et al., 2019) or on stacked galaxies observed up to high redshift (Fornasini et al., 2018, $z<5$). We calculate the combined X-ray luminosity of X-ray binaries in galaxies using the relations of Fornasini et al. (2018).
@@
We model the emission from MBHs by considering the total mass of the binary and the accretion rate recorded in the simulation averaged over 1 Myr around the time where the MBHs merge. This to account that the MBHs can be observed at any point during the time spent in band, and this time is generally of the order of a Myr (see Figure 1).
We model the AGN emission from X-ray to Infrared using the average spectral energy distribution for AGN developed in Shen et al. (2020). We add a correction for dust by estimating the hydrogen column density within a sphere of radius $4\,\rm kpc$, assuming a dust-to-gas mass ratio of $10^{-2}$ and a Milky Way-like extinction curve. For simplicity we fit the hydrogen column density for the MBH population as:
@@
We should mention important caveats. For MBHs with Eddington ratio $\lesssim 10^{-2}-10^{-3}$, the total luminosity should be rescaled to account for radiatively inefficient accretion discs, and the spectral energy distribution differs overall from the radiatively efficient case. We have compared the results with the specific radiatively inefficient solution developed by Nemmen et al. (2014) for a range of masses and accretion rates, and found that rescaling the X-ray to infrared spectral energy distribution by a factor $f_{\rm Edd}$ (Merloni & Heinz, 2008) gives an estimate within a factor about 5 from the Nemmen et al. (2014) solution. Given the large uncertainties, we report the radiatively inefficient sources, which we define as having $f_{\rm Edd}<10^{-3}$, with an optimistic estimate corresponding to the radiatively efficient case, and we discuss the case corresponding to the same multiplied by $f_{\rm Edd}$ to assess the effect of suppressing the radiative efficiency.
@@ 7.2 Electromagnetic properties of MBHs and host galaxies
We compare the brightness of CW candidates and their host galaxies in Figure 14. The host galaxies are shown as black empty pentagons and the AGN as red filled dots for $f_{\rm Edd}>10^{-3}$ and as orange triangles otherwise, optimistically assuming radiatively efficient accretion even at low Eddington ratios. Appendix C compares with radiatively inefficient accretion.
@@
The first point to note is that the large majority of MBHBs have bolometric luminosities below $10^{45}\,\rm erg\,s^{-1}$, the typical dividing line between quasar (brighter sources) and AGN (fainter sources). This is because, although the MBHs are very massive, their typical accretion rates are very low (Figure 11). Only 9% of CW candidates, and 16% of the high probability CW candidates, have $L_{\rm bol}>10^{45}\,\rm erg\,s^{-1}$. More generally, the electromagnetic properties of the CW candidates are very similar to those of high-probability CW candidates. The fraction is about 3% for MBHBs dominating the background (see Appendix C). This is in agreement with Truant et al. (2025). A related, but slightly different question, is the binary fraction in quasars. The fraction of quasars with $L_{\rm bol}>10^{45}\,\rm erg\,s^{-1}$ that host binaries close to coalescence – without a restriction to nHz frequencies – is 7% in our model. Casey-Clyde et al. (2024) also find a low binary fraction in quasars ($\sim 4$%).
@@
In optical and near-IR bands (B, V, and K) the host galaxies are generally bright, with magnitudes brighter than 20, while about half of the AGN have magnitudes fainter than 20. The results are in qualitative agreement with Veronesi et al. (2025), although we do not find a similarly strong dependence of the photometric properties of the binary and its host on the binary mass. Even in this optimistic scenario for the radiative efficiency, none of the MBHBs with $f_{\rm Edd}<10^{-3}$ are brighter than the host galaxies, while for MBHBs with higher accretion rates, 34%, 17% and 26% are brighter than the host in B, V, and K respectively. In radio and X-rays all MBHBs with $f_{\rm Edd}>10^{-3}$ are brighter than the host, while the low accretors can be out-shined by the stellar emission from the host galaxy. If we optimistically do not include a correction for radiative inefficiency about 30% (X-ray) and 80% (radio) can be brighter than the host. Results are qualitatively similar for binaries dominating the background. In Appendix C, we show explicitly the ratio of AGN and host galaxy luminosity for this optimistic case as well as the results when applying a correction for radiatively inefficient sources. Figure 16 highlights how MBHBs with $f_{\rm Edd}<10^{-3}$, which are the majority of the sample, are in the more pessimistic case always fainter than the galaxy emission, even in X-ray and radio. Overall, however, we consider the radio band the best suited for identifying MBHBs with less contamination from the host galaxy, followed by X-rays.
@@
The preponderance of faint AGN out-shined by their host galaxies among CW candidates makes searches for MBHBs harder: variation of the MBHB lightcurve have to be extracted from the brighter non-variable (unless there is a rare event like a supernova) galaxy emission. Electromagnetic searches for MBHBs are therefore often limited to the brightest MBHBs, which are however a minority with respect to the full MBHB population. For instance, Foustoul et al. (2025) searched for MBHBs candidates, looking for continuous periodic modulation in the optical lightcurves (g- and r-bands) of sources coincident with the centers of galaxies. Each candidate found in Foustoul et al. (2025) was shown to be a quasar and the luminosity of the central MBHB candidate outshines the host galaxy. Plotted in Figure 14 are the electromagnetic properties of these MBHB candidates at redshift $0.4<z<1.0$. We show all candidates for which there are K-, V- and B-band photometry (left-hand plots) or X-ray detections/upper limits (right-hand plots). Considering an orbital origin for the optical variability, the separation of these candidates is between 4 and 13 mpc, similar to those in the simulations and making them potential PTA CW sources. We calculated their chirp mass, shown in Figure 14, supposing an equal mass binary and with masses estimated from their SDSS optical spectrum in Rakshit et al. (2020). The bolometric luminosities were estimated using their X-ray detections/upper limits and the bolometric correction computed for type I AGN in Duras et al. (2020). The MBHB candidates have B, V, and K magnitudes slightly fainter than many of the galaxies and AGN in Figure 14 as they are on average at greater distances than the CW candidates in the simulation. The X-ray fluxes are similar to the CW candidates, but again slightly fainter than the AGN due to the greater distances. Accounting for distance (redshift) by considering the X-ray luminosity, the electromagnetic candidates show very similar luminosities as the AGN and indeed all of them are quasars (see also Foustoul et al. (2025)). Calculating the Eddington ratios for the MBHB candidates, we determine ratios between 0.05 and 0.3, indicating high accretion rates, which explains the high bolometric luminosities for these candidates. Overall, the electromagnetic properties of the MBHB candidates are similar to the expected properties of CW from the simulation with the exception that the candidates discovered through optical observations are generally at higher redshifts. This is probably because only $\sim$20% of MBHB are expected to outshine their galaxies in the g- and r-bands, so a larger volume must be searched to identify this rarer sample.

### v2

@@ 6 Properties of MBHs and host galaxies
Based on the analysis presented in the previous sections, here we focus on binaries with chirp mass $10^{8.35}\,{\rm M_{\odot}}$ and $0.05<z<1.2$ as the dominant component of the background (Sect. 4.2.2) and the selected individual CW candidates in population modeling for circular binaries (Sect. 5.2.2) as representative of the population. As a general point of information, a given MBH may appear in the catalog multiple times with multiple redshifts as it merges with different MBHs. This is the case for the most massive MBHs, hosted in very massive galaxies, since they experience numerous mergers in their lifetime. Physically, this does not mean that a given binary in the Universe would be detected multiple times. However, if a simulation is a realistic representation (of a finite volume) of our Universe, the MBH binaries populating a given redshift layer will have similar properties as those merging at the corresponding redshift in the simulation. For a sufficiently large simulation box, which probes the large scale structure of the Universe, this is a reasonable assumption.
@@
First, we consider what it is that makes an MBHB a good CW candidate physically. Besides its position in the sky, near pulsars, and its inclination, which we assumed to be both random in our analysis, the only physical quantities from the simulation are chirp mass and redshift. We consider binaries that make the best CW candidates as those that appear most often in the various Universe realizations. The best CW candidates are not necessarily those that have the largest S/Ns. The largest S/Ns are obtained for the largest chirp masses and the lowest redshifts (luminosity distance). The largest chirp masses have however the shortest residence times in the PTA band (Mingarelli et al., 2017), making it harder to catch them in the act. The lowest redshift sources are rarer because of the small volume probed. As a result, the binaries associated with the very high S/N are also appearing very rarely in the Universe realizations. We therefore consider as ”high probability” CW candidates those that appear multiple times (i.e., at least five times in this work) across our 2 0002\,000 realizations.
@@
In the following, we consider both the simulation box, which is an evolving patch of the Universe that we can analyze at different times, and the simulation lightcone, which mimics what and how an observational survey would “see” the simulation. When we move from the simulation box to the lightcone generated for the simulation (Laigle et al., 2019) a given MBH appears only once, and the cosmic time at which the MBH appears in the lightcone may or may not coincide with the time of an MBH merger. This means that only a small fraction of merging binaries are contained in the lightcone.
@@
In Fig. 9, we compare the properties of mergers dominating the background in the simulation box which has a fixed co-moving volume (teal), in the simulation lightcone (light green) which simulates the source distribution seen by an observer, and the CW candidates (red). In these distributions, we include CW candidates accounting for their multiplicity across Universe realizations, as the multiplicity of a given candidate in the various realizations provides information on how strong a candidate is as it provides information on its likelihood. In a cosmological volume, and in a lightcone that has increasing volume as redshift increases, the background is dominated by the more abundant MBHs with mass $\sim 10^{8.5}\,{\rm M_{\odot}}$, rather than the stronger but rarer GW sources with larger mass. CW candidates are instead mainly found at low redshift and have higher MBH masses; some can also have relatively low masses, down to $\sim 3\times 10^{7}\,{\rm M_{\odot}}$, but they are all at $z<0.1$. The redshift distribution extends to higher redshifts in the lightcone, because of the larger volume probed with increasing redshift, while it is limited to lower redshifts for CW candidates, because the strength of the GW signal decreases with increasing luminosity distance, hence redshift. The degeneracy between luminosity distance (redshift) and chirp mass (see Eqs. 22 and 30) allows for the presence of CW candidates up to $z=0.4$, provided that they have chirp mass $\mathcal{M}_{\rm c}\sim 10^{9.5}\,\,{\rm M_{\odot}}$. The distribution of the masses of the host galaxies reflects the $\mathcal{M}_{\rm c}$ and $z$ distributions for the background, while for CW candidates it is dominated by massive galaxies, in line with the larger MBH masses with respect to the background. These results are in agreement with previous and concurrent investigations (e.g., Sesana et al., 2009; Cella et al., 2025; Truant et al., 2025).
@@
We examine the relation between binaries and environment in Fig. 10. Here, $\mathcal{M}_{\rm c}$ is shown as a function of the halo mass of the host. We note some “undermassive binaries,” with very low chirp mass with respect to the host galaxy/halo: these are cases where the mass ratio in the binary is very small, and therefore the chirp mass differs significantly from the binary mass, which is what scales with host properties in the simulation (shown in Volonteri et al., 2020). The binaries dominating the background in the simulation box are primarily hosted in galaxies and halos that are more massive than the Milky Way, while CW candidates are, for the most part, hosted in group-like main halos (red squares) with $M_{\rm halo}>10^{13}\,{\rm M_{\odot}}$. CW candidates that are hosted in lower mass halos are all located in subhalos (orange crosses), meaning that the subhalo has been affected by stripping and it is part of a larger halo [^10] The plume of binaries in the background with halo mass $10^{12}\,{\rm M_{\odot}}$ is also caused by subhalos.; the high chirp mass with respect to halo mass is explained by stripping of the halo, which has therefore lost part of its mass (see the discussion in Volonteri et al., 2016, for the general MBH population). This implies that CW candidates are best looked for in galaxy groups and clusters, even more so than for the general population of MBH binaries (Rosado and Sesana, 2014; Izquierdo-Villalba et al., 2023; Saeedzadeh et al., 2024). Using groups/clusters as priors for where to look for CW candidates in the Bayesian search of PTA data could be therefore a possible strategy.
@@
Furthermore, the MBHs in the central galaxies of overdensities experience a large number of mergers, up to $\sim 20-25$ for the MBHBs that dominate the nHz background. This is a manifestation of the significant fraction of MBH mass gained through MBH mergers for these galaxies (Volonteri and Ciotti, 2013; Dubois et al., 2014b; Kulier et al., 2015). This occurs because of a combination of enhanced galaxy mergers and decreased gas availability for accretion for high-mass MBHs hosted in massive central galaxies of groups and clusters.
@@
In Fig. 11, we compare the Eddington ratios and star formation rate (SFR) of background sources in the box vs CW candidates. We average SFR over 100 Myr since most star formation diagnostics probe SFR on such timescales and we average $f_{\rm Edd}$ over 1 Myr to smooth the variability. Also in this case we include CW candidates accounting for their multiplicity across Universe realizations. The Eddington ratio of the binaries is generally low, $10^{-2}$, typical for MBHs of this mass and redshift in the simulation as well as in observations (Volonteri et al., 2016). Accretion is even lower for CW candidates, since they are at very low redshift, in very massive galaxies. The same considerations apply to SFR, in agreement with previous results which find that the host has low SFR or specific SFR (Saeedzadeh et al., 2024; Truant et al., 2025; Cella et al., 2025). Since we are considering delayed mergers rather than numerical mergers, the time elapsed from galaxy merger to binary observation is sufficiently long that any galaxy merger-driven increase in accretion rate or SFR has since decayed (Dong-Páez et al., 2023).
@@
Examples of hosts of CW candidates in false gri colors are shown in Fig. 12 and Fig. 13. The MBHB is located at the center of the image. The first three images (Fig. 12 and Fig. 13 left) show two low-redshift cases consistently appearing in the realizations, while Fig. 13, right, shows a rarer high-redshift ($z=0.42$) case. The images highlight that CW candidates are preferentially found in the center of galaxy groups and clusters, in agreement with the results shown in Fig. 10.
@@ 7 Electromagnetic emission from galaxies and MBHs
Figure 14: Electromagnetic properties of MBHBs and their host galaxies for CW candidates. The left panels show B, V, and K magnitudes; the right panels show X-ray flux at [2-10] keV, radio flux at 2 GHz and bolometric luminosity (the horizontal line marks $L_{\rm bol}=10^{45}\,\rm{erg}\,{\rm s^{-1}}$), as a typical value for quasar luminosities. For MBHs with $f_{\rm Edd}<10^{-3}$ we show an optimistic upper limit to the brightness ignoring the correction to radiative efficiency.
@@ 7.1 Methodology
Galaxies magnitudes are computed as described in Laigle et al. (2019). In brief, we compute the galaxy rest-frame spectral energy distribution by summing up the contribution of all stellar particles (knowing their ages and metallicities), assumed to behave as single stellar populations. We use the stellar population synthesis model of Bruzual and Charlot (2003) assuming a Chabrier (2003) initial mass function. The dust column density and optical depth along the line of sight of each star particle is computed assuming that 30% of metals in the galaxy gas distribution are locked into dust grains. We use the $R_{\rm V}=3.1$ Milky Way dust grain model by Weingartner and Draine (2001). Apparent total magnitudes are computed on the redshifted spectra, using the $\rm B$, $\rm V$ and $\rm K_{\rm s}$ filters.
In X-rays, the galaxy emission is dominated by X-ray binaries composed of a compact object accreting from a companion star. Empirical models allow for the combined effect of low-mass X-ray binaries to be estimated, where the companion is a low-mass star and the emission depends on the total galaxy mass, and high-mass X-ray binaries, where the companion is a high-mass star and the emission depends on the SFR. These models are based either on individual observations of nearby galaxies (e.g., $50\rm\,Mpc$, Lehmer et al., 2019) or on stacked galaxies observed up to high redshift (Fornasini et al., 2018, $z<5$). We calculate the combined X-ray luminosity of X-ray binaries in galaxies using the relations of Fornasini et al. (2018).
@@
We model the emission from MBHs by considering the total mass of the binary and the accretion rate recorded in the simulation averaged over 1 Myr around the time when the MBHs merge. This is to account for that the MBHs can be observed at any point during the time spent in band, and this time is generally of the order of a Myr (see Fig. 1).
We model the AGN emission from X-ray to infrared (IR) using the average spectral energy distribution for AGN developed in Shen et al. (2020). We add a correction for dust by estimating the hydrogen column density within a sphere of radius $4\,\rm kpc$, assuming a dust-to-gas mass ratio of $10^{-2}$ and a Milky Way-like extinction curve. For simplicity we fit the hydrogen column density for the MBH population as
@@
We should mention important caveats. For MBHs with Eddington ratio $\lesssim 10^{-2}-10^{-3}$, the total luminosity should be rescaled to account for radiatively inefficient accretion discs, and the spectral energy distribution differs overall from the radiatively efficient case. We compared the results with the specific radiatively inefficient solution developed by Nemmen et al. (2014) for a range of masses and accretion rates, and found that rescaling the X-ray to IR spectral energy distribution by a factor $f_{\rm Edd}$ (Merloni and Heinz, 2008) gives an estimate within a factor of about 5 from the Nemmen et al. (2014) solution. Given the large uncertainties, we report the radiatively inefficient sources, which we define as having $f_{\rm Edd}<10^{-3}$, with an optimistic estimate corresponding to the radiatively efficient case, and we discuss the case corresponding to the same multiplied by $f_{\rm Edd}$ to assess the effect of suppressing the radiative efficiency.
@@ 7.2 Electromagnetic properties of MBHs and host galaxies
We compare the brightness of CW candidates and their host galaxies in Fig. 14. The host galaxies are shown as black empty pentagons and the AGN as red filled dots for $f_{\rm Edd}>10^{-3}$ and as orange triangles otherwise, optimistically assuming radiatively efficient accretion even at low Eddington ratios. Appendix C compares with radiatively inefficient accretion.
@@
The first point to note is that the large majority of MBHBs have bolometric luminosities below $10^{45}\,\rm erg\,s^{-1}$, the typical dividing line between quasars (brighter sources) and AGN (fainter sources). This is because, although the MBHs are very massive, their typical accretion rates are very low (Fig. 11). Only 9% of CW candidates, and 16% of the high probability CW candidates, have $L_{\rm bol}>10^{45}\,\rm erg\,s^{-1}$. More generally, the electromagnetic properties of the CW candidates are very similar to those of high-probability CW candidates. The fraction is about 3% for MBHBs dominating the background (see Appendix C). This is in agreement with Truant et al. (2025). A related, but slightly different question, is the binary fraction in quasars. The fraction of quasars with $L_{\rm bol}>10^{45}\,\rm erg\,s^{-1}$ that host binaries close to coalescence – without a restriction to nHz frequencies – is 7% in our model. Casey-Clyde et al. (2025) also find a low binary fraction in quasars ($\sim 4$%).
@@
In the optical and near-IR (NIR) bands (B, V, and K) the host galaxies are generally bright, with magnitudes brighter than 20, while about half of the AGN have magnitudes fainter than 20. The results are in qualitative agreement with Veronesi et al. (2025), although we do not find a similarly strong dependence of the photometric properties of the binary and its host on the binary mass. Even in this optimistic scenario for the radiative efficiency, none of the MBHBs with $f_{\rm Edd}<10^{-3}$ are brighter than the host galaxies, while for MBHBs with higher accretion rates, 34%, 17% and 26% are brighter than the host in B, V, and K respectively. In radio and X-rays, all MBHBs with $f_{\rm Edd}>10^{-3}$ are brighter than the host, while the low accretors can be outshined by the stellar emission from the host galaxy. If we optimistically do not include a correction for radiative inefficiency, about 30% (X-ray) and 80% (radio) can be brighter than the host. Results are qualitatively similar for binaries dominating the background. In Appendix C, we explicitly show the ratio of AGN and host galaxy luminosity for this optimistic case as well as the results when applying a correction for radiatively inefficient sources. Fig. 16 highlights how MBHBs with $f_{\rm Edd}<10^{-3}$, which are the majority of the sample, are in the more pessimistic case always fainter than the galaxy emission, even in X-ray and radio. Overall, however, we consider the radio band the best suited for identifying MBHBs with less contamination from the host galaxy, followed by X-rays.
@@
The preponderance of faint AGN outshined by their host galaxies among CW candidates makes searches for MBHBs harder: variations of the MBHB light curve have to be extracted from the brighter nonvariable (unless there is a rare event like a supernova) galaxy emission. Electromagnetic searches for MBHBs are therefore often limited to the brightest MBHBs, which are however a minority with respect to the full MBHB population. For instance, Foustoul et al. (2025) searched for MBHB candidates, looking for continuous periodic modulation in the optical light curves (g- and r-bands) of sources coincident with the centers of galaxies. Each candidate found in Foustoul et al. (2025) was shown to be a quasar and the luminosity of the central MBHB candidate outshines the host galaxy. Plotted in Fig. 14 are the electromagnetic properties of these MBHB candidates at redshift $0.4<z<1.0$. We show all candidates for which there are K-, V- and B-band photometry (left-hand plots) or X-ray detections/upper limits (right-hand plots). Considering an orbital origin for the optical variability, the separation of these candidates is between 4 and 13 mpc, similar to those in the simulations and making them potential PTA CW sources. We calculated their chirp mass, shown in Fig. 14, supposing an equal mass binary and with masses estimated from their SDSS optical spectrum in Rakshit et al. (2020). The bolometric luminosities were estimated using their X-ray detections/upper limits and the bolometric correction computed for type I AGN in Duras et al. (2020). The MBHB candidates have B, V, and K magnitudes slightly fainter than many of the galaxies and AGN in Fig. 14 as they are on average at greater distances than the CW candidates in the simulation. The X-ray fluxes are similar to the CW candidates, but again slightly fainter than the AGN due to the greater distances. Accounting for distance (redshift) by considering the X-ray luminosity, the electromagnetic candidates show very similar luminosities as the AGN and indeed all of them are quasars (see also Foustoul et al. 2025). Calculating the Eddington ratios for the MBHB candidates, we determined ratio values between 0.05 and 0.3, indicating high accretion rates that would explain the high bolometric luminosities for these candidates. Overall, the electromagnetic properties of the MBHB candidates are similar to the expected properties of CW from the simulation with the exception that the candidates discovered through optical observations are generally at higher redshifts. This is probably because only $\sim$20% of MBHBs are expected to outshine their galaxies in the g- and r-bands, so a larger volume must be searched to identify this rarer sample.